Five gallons of a Php 120 per gallon thinner is to be mixed with a paint costing Php 760 per gallon to make a paint that can be sold for Php 560 per gallon. How much paint will be produced?
Answer by richwmiller(17219)

7.6*x+1.2*5=5.6(5+x)
7.6x-5.6x=28-6
2x=22
x=11
5+11=16 gallons total